OpenStudy (anonymous):

f(1) = 2 and f(2) = 3, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______

OpenStudy (anonymous):

I think it's 10, but I'm not sure.

OpenStudy (anonymous):

how do you solve it? bc im lost even how to get the answer :/

OpenStudy (anonymous):

f(n) = f(1) + f(2) + f(n-1), right? f(5) = 2 + 3 + f(4) Assuming that f(x) = x+1, f(4) would be 5 f(5) = 2 + 3 + 5 f(5) = 10

OpenStudy (anonymous):

OH ok thank you so much!

OpenStudy (loser66):

for n>2 , apply the formula so, let take n=3, f(3) = f(1) +f(2) + f(3-1) = f(1) + 2f(2) = 1+ 2*2= 5 n=4, f(4) = f(1) + f(2) +f(4-1) = f(1)+f(2) + f(3) = 1+2 +5 = 8 n=5 , f(5) = f(1) +f(2) +f(5-1) = f(1)+f(2) + f(4) = 1+2+8 = 11

OpenStudy (loser66):

Does it make sense?

OpenStudy (anonymous):

yes thanks

OpenStudy (whpalmer4):

uh @loser66 your arithmetic isn't quite right... \[f(3) = f(1) + f(2) + f(3-1) = f(1)+ f(2) + f(2) = f(1) + 2f(2) \]\[\qquad = 2+2*3 = 8\]

OpenStudy (loser66):

yes, hahaha... thank you @whpalmer4

OpenStudy (anonymous):

so.... now im lost....

OpenStudy (loser66):

guide him, please, @whpalmer4

OpenStudy (whpalmer4):

then \(f(4) = f(1)+f(2) + f(4-1) = 2+3 + 8 = \)

OpenStudy (loser66):

I am not good at teaching. :(

OpenStudy (anonymous):

that = 13 so now what? or is 13 the answer?

OpenStudy (whpalmer4):

so \(f(4) = 13\) but we need to find \(f(5)\) can you tell me the expression for \(f(5)\) from the formula above?

OpenStudy (anonymous):

so then would it be f(5)=f(1)+f(2)+f(5-1)?

OpenStudy (whpalmer4):

exactly! and we know \(f(5-1) = f(4) = 13\)

OpenStudy (anonymous):

so then 2+3+13=18 so the answer is 18?

OpenStudy (whpalmer4):

yes, that's my belief...

OpenStudy (anonymous):

ok, give me a couple of minutes and ill be able to tell you if that's right or not lol

OpenStudy (whpalmer4):

First 10 values: 2, 3, 8, 13, 18, 23, 28, 33, 38, 43

OpenStudy (whpalmer4):

actually it's just a simple arithmetic sequence after the first 2 values, because we always add 2+3 to the previous term...

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

i submitted 18 and its apparently wrong....

OpenStudy (whpalmer4):

are you positive you have shown us the problem exactly as stated? Those values I gave you are the values of the recursive sequence you described.

OpenStudy (whpalmer4):

you had f(2) = 3 in the problem statement here corrected sequence is 2, 5, 12, 19, 26, 33, 40, 47, 54, 61

OpenStudy (anonymous):

sorry... so the answer is 26?

OpenStudy (whpalmer4):

yes, the terms in my list are f(1), f(2), f(3), f(4), f(5), etc. \[f(5) = f(1) + f(2) + f(4) = f(1) + f(2) + (f(1)+f(2)+(f(1)+f(2)+f(2))) \]\[\qquad= 2+5+(2+5+(2+5+5)) =26\]

OpenStudy (anonymous):

ok thx let me try that one

OpenStudy (whpalmer4):

gets a little confusing typing it like that :-) maybe better would have been \[f(5) = f(1)+f(2) + f(4)\]\[f(4) = f(1)+f(2) + f(3)\]\[f(3) = f(1)+f(2)+f(2)\]\[f(2) = 5\]\[f(1) =2\]\[f(5) = 2+5+(2+5+(2+5+5)) = \] this way does make it apparent that now we have an arithmetic sequence with a common difference of 7 instead of 5, because each additional term adds 2+5=7

OpenStudy (anonymous):

ok so im being asked for f(1) = 2 and f(2) = 3, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______ that would be 18?

OpenStudy (whpalmer4):

is this a different problem? in the screen capture you sent, it was f(2) = 5...

OpenStudy (anonymous):

yes it is

OpenStudy (whpalmer4):

Okay, then yes, if you work it out the same way, you get \(f(5) = 18\) just as we did when we were doing the "wrong" problem :-)

OpenStudy (anonymous):

lol so it is 18

OpenStudy (whpalmer4):

probably best way for you to write it would be\[f(1) = 2\]\[f(2) = 3\]\[f(3) = f(1)+f(2)+f(2) = 2+3 +3 = 8\]\[f(4) = f(1)+f(2) + f(4-1) = f(1)+f(2) + f(3) = 2+3+8 = 13\]\[f(5) = f(1)+f(2) + f(5-1) = f(1)+f(2) + f(4) = 2+3+13 = 18\] rather than trying to do it recursively like I did

OpenStudy (whpalmer4):

better middle line: \[f(3) = f(1)+f(2) + f(3-1) = f(1)+f(2)+f(2) = 2+3+3=8\]

OpenStudy (whpalmer4):

baby steps, baby steps :-)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

WHOOOOOO!!!!!!!!! that was the right answer thank you soooooooooooooo much!!

OpenStudy (whpalmer4):

Great!